# Tally: Non-Intrusive Performance Isolation for Concurrent Deep Learning Workloads

### v1

@@ Abstract.
GPU underutilization is a significant concern in many production deep learning clusters, leading to prolonged job queues and increased operational expenses. A promising solution to this inefficiency is GPU sharing, which improves resource utilization by allowing multiple workloads to execute concurrently on a single GPU. However, the practical deployment of GPU sharing in production settings faces critical obstacles due to the limitations of existing mechanisms, such as high integration costs, inadequate performance isolation, and limited application compatibility. To address these issues, we introduce Tally, a non-intrusive GPU sharing mechanism that provides robust performance isolation and comprehensive workload compatibility. Tally operates as a virtualization layer between applications and GPUs, transparently orchestrating the device execution of concurrent workloads. The key to Tally’s robust performance isolation capability lies in its fine-grained thread-block level GPU kernel scheduling strategy, which allows the system to effectively mitigate interference caused by workload co-execution. Our evaluation, conducted on a diverse set of workload combinations, demonstrates that Tally on average incurs a mere $7.2\%$ overhead on the $99^{th}$-percentile latency of high-priority inference tasks when executed concurrently with best-effort training workloads compared to $188.9\%$ overhead exhibited by the state-of-the-art GPU sharing systems like TGS, while achieving over $80\%$ of TGS’s system throughput.
Part of this work was done while the author was a student at the University of Toronto.
@@ 1. INTRODUCTION
Deep learning (DL) has demonstrated astonishing performance across a wide range of applications, including image recognition (krizhevsky2012, ; kaiming2016, ; szegedy2015, ), machine translation (devlin-etal-2019-bert, ; bahdanau2015neural, ; sutskever2014, ; lewis-etal-2020-bart, ), recommendation systems (Zhang2020, ; alashkar2017examples, ), and autonomous driving  (Cordts2016, ; yurtsever2020survey, ). Recent advances in generative AI (rombach2022high, ; brown2020language, ), exemplified by the development of large language models (LLMs) like ChatGPT (openai2024chatgpt, ), have further elevated the influence of DL to unprecedented levels. To facilitate the research and deployment of DL-based applications, many organizations (shukla2022singularity, ; Weng2022mlaas, ; meta2024, ) are investing in building computing infrastructures with large fleets of GPU devices. For example, Meta is building two data center-scale clusters for AI, each equipped with hundreds of thousands of NVIDIA H100 GPUs with an estimated cost as high as half a billion dollars (meta2024, ; Eadline_2023, ). As the demand for computational resources continues to escalate, efficient utilization of GPUs becomes paramount to mitigate costs and ensure scalability for AI applications.
@@
Nevertheless, studies have noted significant GPU underutilization in large-scale DL clusters. For example, a recent study by Alibaba (Weng2022mlaas, ) revealed that the median GPU utilization in their production DL cluster can go as low as $4.2\%$. This inefficiency results from the need for production clusters to accommodate a diverse range of workloads, many of which are unable to fully leverage the capacity of the GPUs. For instance, training processes can potentially underutilize GPUs due to bottlenecks in CPU execution such as stalls caused by data fetching and preprocessing (Audibert2023, ). The utilization of online inference tasks, on the other hand, is intrinsically tied to real-time request rates, which are subject to substantial fluctuations causing unpredictable GPU idleness (li2023-alpaserver, ). For instance, Microsoft’s production trace (maf2, ) has revealed significant burstiness in user requests, with frequent spikes in demand of up to 50×50$\times$50 × the average. This also makes usage prediction and resource allocation challenging for inference workloads, resulting in resource over-provisioning to ensure high quality of service (MArk2019, ; li2023-alpaserver, ), which exacerbates the underutilization problem.
@@
To enhance the efficiency of DL clusters, recent studies (xiao2020, ; salus2020, ; lucid, ; narayanan2020heterogeneity, ) suggest GPU sharing as a promising solution. This approach aims to enhance hardware utilization by allowing multiple low-GPU tasks to share a single GPU’s resources. The same Alibaba study conducted a simulation on their production traces and demonstrated that an effective GPU sharing mechanism could potentially reduce the amount of GPU resources required by the cluster by $50\%$ on average and up to $73\%$ during peak hours. These findings underscore the potential of GPU sharing in mitigating the GPU underutilization problem and significantly improving the overall efficiency of large-scale DL clusters.
@@
Despite these promising results, GPU sharing has not yet achieved widespread adoption in production. One major obstacle is the lack of performance isolation mechanisms in most current GPU sharing systems. Many workloads within DL clusters are governed by service-level agreements (SLAs), which impose stringent performance requirements, usually on the order of milliseconds (clipper2017, ). These workloads are commonly referred to as high-priority or latency-critical tasks. By contrast, workloads that can tolerate extended processing times are known as low-priority or best-effort tasks. While it seems logical to pack best-efforts tasks along with high-priority tasks to utilize the GPU idle time caused by the latter, this approach can cause significant interference to the performance of the high-priority workloads, potentially violating their SLAs. Our experiments on state-of-the-art GPU sharing mechanisms reveal that such performance interference can cause up to a 20×20$\times$20 × slowdown in tail latency (see Section 5), making them impractical for use in practice.
@@
Another hindrance to the broader adoption of GPU sharing lies in the intrusive nature of many current sharing solutions  (salus2020, ; xiao2020, ; case2022, ; effisha2017, ; slate2019, ; tacker2022, ). Using such solutions require extensive modifications to user-level code or the underlying DL frameworks (e.g., PyTorch  (paszke2019pytorch, ), TensorFlow (tensorflow2016, )). These solutions not only greatly limit the flexibility of users in developing their workloads but also impose a substantial burden on the users and the cluster providers for workload adaptation and framework upgrades, leading to significant integration and maintenance costs (lucid, ). Moreover, many of these solutions require the workloads to adhere to specific characteristics, such as ensuring GPU kernels used by the workload are always idempotent (reef2022, ). These constraints often only apply narrowly in production clusters that host a diverse range of workloads with varying characteristics.
@@
In light of these observations, we argue that for practical deployment in production DL clusters, a GPU sharing system must meet the following three criteria. (1) Non-intrusive to seamlessly integrate into existing clusters and DL workloads, (2) provide performance isolation guarantees to ensure that the service-level requirements of high-priority tasks are maintained effectively during shared execution, and (3) be generalizable across a diverse range of DL applications. Building on these objectives, we propose Tally, a novel non-intrusive GPU sharing system for large-scale production DL clusters that ensures performance isolation of latency-critical high-priority tasks when co-located with best effort tasks, while maximizing hardware efficiency and overall system throughput. We design Tally as a transparent virtualization layer between the application-level and the GPU. It intercepts GPU kernel launch API calls and employs a task-agnostic scheduling algorithm to prioritize high-priority kernels, while opportunistically scheduling best-effort kernels during GPU idle periods.
@@
To efficiently co-locate high-priority and best-effort tasks on a GPU with performance isolation guarantees, it is essential to schedule best-effort kernels in a controlled manner when the GPU is idle and promptly switch to high-priority kernels as required. However, modern GPUs lack kernel scheduling primitives to terminate a running kernel and only allow context switching between kernels at the granularity of entire kernels (reef2022, ; GILMAN2021102234, ). We observe that this limitation introduces significant overhead on high-priority tasks, especially since many best-effort DL workloads involve long-running kernels. To overcome this challenge and enable finer-grained kernel scheduling, Tally uses two software-level kernel scheduling primitives: slicing and preemption. Slicing divides large kernels into smaller segments, while preemption allows for the interruption of active kernels. Our key insight is that these two scheduling primitives can be implemented non-intrusively and in a task-agnostic manner through a series of GPU kernel transformation passes, leveraging the GPU programming model which organizes kernels as a collection of independent thread blocks.
@@
While slicing and preemption are effective mechanisms to manage co-execution interference, each has its trade-offs. Kernel slicing is a lightweight transformation but can incur higher overhead due to multiple kernel launches. In contrast, kernel preemption requires only a single kernel launch but involves a more complex transformation with potential synchronization overhead. Due to these differences, the optimal choice of scheduling primitives and configuration parameters (e.g., the degree of slicing) varies across different kernels. Towards this end, Tally uses a profile-guided resource provisioning strategy in its scheduling algorithm that transparently profiles the performance of best-effort kernels on-the-fly across various scheduling configurations to select the best option to meet the latency requirements of the high-priority task. This approach ensures that interference from best-effort tasks remains within acceptable limits while striving for optimal overall system throughput.
@@
To evaluate Tally’s ability to provide effective performance isolation, we prepare a comprehensive benchmark suite consisting of six training and six inference workloads that cover a wide range of popular DL models and frameworks (details shown in table 2). We evaluate the co-located performance of latency-critical inference tasks and best-effort training tasks by simulating the production traffic patterns for the inference tasks using the publicly available Microsoft Azure Function Trace 2021 (MAF2) dataset (maf2, ). Across all the combinations, Tally incurs a mere $7.2\%$ overhead on average in the $99^{th}$-percentile latency of inference tasks, compared to $252.3\%$, $345.0\%$, $195.5\%$, and $188.9\%$ exhibited by state-of-the-art non-intrusive GPU sharing solutions, Time-Slicing, MPS, MPS-Priority, and TGS respectively. At the same time, Tally is able to achieve $80\%$ of the system throughput attained by TGS, demonstrating the effectiveness of Tally in maximizing the GPU efficiency while providing strong performance isolation guarantees.
@@
  • We introduce a novel approach to achieve fine-grained GPU kernel scheduling in a non-intrusive and task-agnostic manner using kernel slicing and preemption primitives. We further propose a priority-aware kernel scheduler that dynamically selects the appropriate scheduling primitives to optimize GPU throughput while preserving performance isolation guarantees.
  • We evaluate Tally on a diverse range of inference and training workloads and show that Tally can achieve system throughput on-par with state-of-the-art GPU sharing solutions while providing strong performance isolation guarantees.
@@ 2. BACKGROUND
Modern large-scale DL clusters are typically shared among many users within an organization and are designed to support a diverse range of workloads (xiao2020, ; lucid, ), encompassing training, fine-tuning, and inference tasks for a variety of model architectures, such as convolutional neural networks (CNNs), transformer-based large language models (LLMs), graph neural networks (GNNs), and reinforcement learning (RL) models. These tasks are often implemented using various mainstream DL frameworks such as PyTorch, TensorFlow, and JAX, as well as proprietary softwares (Weng2022mlaas, ). Given the diversity and compute-intensive nature of DL workloads, general-purpose GPUs have become the dominant accelerator in DL clusters due to their powerful programming model and immense parallel computing capabilities.
@@
GPU architecture and programming model. While there are variations among vendors (e.g, NVIDIA and AMD), modern GPUs used in DL clusters fundamentally follow similar architecture and programming model. GPUs are equipped with thousands of parallel processing units called Streaming Multiprocessors (SMs). Kernels, the computational units executed on GPUs, are programmed using specialized languages like CUDA (cuda2024, ). GPU kernels follow a hierarchical thread structure consisting of threads, warps, thread blocks, and grids. Threads within a kernel identify themselves using built-in variables such as threadIdx and blockIdx to determine their task assignments. Warps are groups of threads that execute in SIMT (Single Instruction Multiple Thread) fashion. All the warps within a thread-block are scheduled to a single SM and can cooperate using fast on-chip shared memory. Additionally, threads within a block can synchronize their execution to coordinate memory accesses using synchronization primitives. Finally, thread blocks within a grid can be scheduled and executed independently in any order, either sequentially or in parallel (cuda2024, ). This independent scheduling capability allows thread blocks to scale effectively across multiple SMs.
@@
GPU sharing systems. The diverse nature of DL workloads, coupled with their inherent inefficiencies, has raised significant concerns about GPU underutilization in many production DL clusters (hu2021characterization, ; xiao2020, ). To mitigate this problem, researchers have proposed numerous cluster-level optimizations (xiao2018, ; xiao2020, ; wang2021horizontally, ; salus2020, ), among which GPU sharing is a promising approach to improve utilization by consolidating multiple low-utilization jobs onto the same GPU. From an implementation perspective, current GPU sharing solutions can be broadly categorized into two levels: application and runtime. Application-level solutions propose a new programming interface for writing efficient DL workloads that can be executed concurrently. For example, Salus (salus2020, ) exposes a set of scheduling primitives for defining the iteration boundaries of a training job, enabling the scheduler to manage different training workloads at the iteration-level granularity. Similarly, HFTA (wang2021horizontally, ) provides APIs that allow developers to write multiple hyper-parameter tuning jobs as a single application, which can then be efficiently executed on a single GPU.
@@
Unlike application-level solutions that necessitate modifications to user-level applications, runtime-level GPU sharing solutions operate transparently by integrating directly into either DL frameworks (xiao2020, ) (e.g., PyTorch (paszke2019pytorch, ) or TensorFlow (tensorflow2016, )), container orchestration systems (tgs2023, ) (e.g., Kubernetes (kubernetes, )), or the GPU driver (reef2022, ). These systems typically perform resource sharing at a more fine-grained level. For instance, TGS (tgs2023, ) is designed to be deployed on GPU containers in Kubernetes and schedules workloads at the GPU kernel level, while REEF (reef2022, ) extends GPU drivers to offer more fine-grained thread-level scheduling capabilities. The transparent nature of many these solutions allows for adoption across a wider range of applications with minimal effort. However, designing these solutions is often challenging due to the lack of application-level context.

### v3

@@ Abstract.
GPU underutilization is a significant concern in many production deep learning clusters, leading to prolonged job queues and increased operational expenses. A promising solution to this inefficiency is GPU sharing, which improves resource utilization by allowing multiple workloads to execute concurrently on a single GPU. However, deploying GPU sharing in production settings faces critical obstacles due to the limitations of existing mechanisms, including high integration costs, inadequate performance isolation, and limited application compatibility. To address these issues, we introduce Tally, a non-intrusive GPU sharing mechanism that provides robust performance isolation and comprehensive workload compatibility. The key to Tally’s robust performance isolation capability lies in its fine-grained thread-block-level GPU kernel scheduling strategy, which allows the system to effectively mitigate interference caused by workload co-execution. We evaluate Tally on a diverse range of workloads and show that it incurs an average overhead of only $7.2\%$ on the $99^{th}$-percentile latency of high-priority inference tasks when executed concurrently with best-effort training workloads, compared to $188.9\%$ overhead exhibited by the state-of-the-art GPU sharing systems like TGS, while achieving over $80\%$ of TGS’s system throughput.
GPU sharing, performance isolation, cloud infrastructure, deep learning, systems for machine learning
journalyear: 2025copyright: rightsretainedconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1; March 30–April 3, 2025; Rotterdam, Netherlandsbooktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1 (ASPLOS ’25), March 30–April 3, 2025, Rotterdam, Netherlandsdoi: 10.1145/3669940.3707282isbn: 979-8-4007-0698-1/25/03ccs: Computing methodologies Concurrent algorithmsccs: Computing methodologies Artificial intelligenceccs: Computer systems organization Parallel architectures
@@ 1. INTRODUCTION
Deep learning (DL) has demonstrated astonishing performance across a wide range of applications, including image recognition (krizhevsky2012, ; kaiming2016, ; szegedy2015, ), machine translation (devlin-etal-2019-bert, ; bahdanau2015neural, ; sutskever2014, ; lewis-etal-2020-bart, ), recommendation systems (Zhang2020, ; alashkar2017examples, ), and autonomous driving  (Cordts2016, ; yurtsever2020survey, ). Recent advances in generative AI (rombach2022high, ; brown2020language, ), exemplified by the development of large language models (LLMs) such as ChatGPT (openai2024chatgpt, ), have further elevated the influence of DL to unprecedented levels. To facilitate the research and deployment of DL-based applications, many organizations (shukla2022singularity, ; Weng2022mlaas, ; meta2024, ) are investing in building computing infrastructures with large fleets of GPU devices. For example, Meta has built two data center-scale clusters for AI, each equipped with hundreds of thousands of NVIDIA H100 GPUs with an estimated cost as high as half a billion dollars (meta2024, ; Eadline_2023, ). As the demand for computational resources continues to escalate, efficient utilization of GPUs becomes paramount to mitigate costs and ensure scalability for AI applications.
@@
Nevertheless, studies have noted significant GPU underutilization in large-scale DL clusters. For example, a recent study by Alibaba revealed that the median GPU utilization in their production DL cluster was only $4.2\%$ (Weng2022mlaas, ). This inefficiency arises from the need for production clusters to accommodate a diverse range of workloads, many of which are unable to fully utilize the capacity of GPUs. Training workloads, for instance, may underutilize GPUs due to bottlenecks in CPU execution such as stalls caused by data fetching and preprocessing (Audibert2023, ). On the other hand, the utilization of online inference tasks is intrinsically tied to real-time request rates, which are subject to substantial fluctuations causing unpredictable GPU idleness (li2023-alpaserver, ). For example, Microsoft’s production trace (maf2, ) has revealed significant burstiness in user requests, with frequent spikes in demand up to 50×50$\times$50 × the average. This makes usage prediction and resource allocation challenging for inference workloads, resulting in resource over-provisioning to ensure high quality of service (MArk2019, ; li2023-alpaserver, ), which exacerbates the underutilization problem.
@@
To enhance the efficiency of DL clusters, recent studies (xiao2020, ; salus2020, ; lucid, ; narayanan2020heterogeneity, ) suggest GPU sharing as a promising solution. This approach aims to enhance hardware utilization by allowing multiple low-utilization tasks to share a single GPU’s resources. The same Alibaba study conducted a simulation on their production traces and demonstrated that an effective GPU sharing mechanism could reduce the amount of GPU resources required by the cluster by $50\%$ on average, with reductions reaching up to $73\%$ during peak demand periods. These findings underscore the potential of GPU sharing to mitigate GPU underutilization and significantly improve the efficiency of large-scale DL clusters.
@@
Despite these promising results, GPU sharing has not yet achieved widespread adoption in production. One major obstacle is the lack of robust performance isolation mechanisms in most current GPU sharing systems. Many workloads in DL clusters are governed by service-level agreements (SLAs) with strict performance requirements, often measured in milliseconds (clipper2017, ). These workloads are commonly referred to as high-priority or latency-critical tasks. In contrast, workloads that can tolerate extended processing times are known as low-priority or best-effort tasks. While it seems logical to pack best-efforts tasks along with high-priority tasks to utilize potential GPU idle time, this approach can cause significant interference to the performance of high-priority workloads, violating their SLAs. Our experiments on state-of-the-art GPU sharing mechanisms reveal that such interference can cause up to a 20×20$\times$20 × slowdown in tail latency (see Section 5), making them impractical for production.
@@
Another hindrance to the broader adoption of GPU sharing lies in the intrusive nature of many existing solutions  (salus2020, ; xiao2020, ; case2022, ; effisha2017, ; slate2019, ; tacker2022, ). These systems often necessitate extensive modifications to user code or the underlying DL frameworks (e.g., PyTorch  (paszke2019pytorch, ), TensorFlow (tensorflow2016, )). Such requirements not only greatly limit users’ flexibility in developing workloads but also impose substantial burdens on users and cluster providers for workload adaptation and framework upgrades, leading to considerable integration and maintenance costs (lucid, ). Moreover, many of these solutions require the workloads to adhere to specific characteristics, such as ensuring GPU kernels used by the workload are always idempotent (reef2022, ). These constraints significantly limit their applicability in production clusters, which typically host a diverse range of workloads with heterogeneous characteristics.
@@
In light of these observations, we argue that for practical deployment in production DL clusters, a GPU sharing system must meet the following three criteria: (1) it must be non-intrusive to seamlessly integrate into existing clusters and DL workloads, (2) provide performance isolation guarantees to ensure that the SLAs of high-priority tasks are effectively maintained during shared execution, and (3) be generalizable across a diverse range of DL applications. Building on these objectives, we propose Tally, a novel non-intrusive GPU sharing system for large-scale production DL clusters that ensures robust performance isolation for high-priority tasks when co-located with best-effort workloads, while maximizing hardware efficiency and overall system throughput. We design Tally as a transparent virtualization layer between the application level and the GPU. It intercepts GPU kernel launch API calls and employs a task-agnostic scheduling strategy that prioritizes the execution of high-priority kernels while opportunistically scheduling best-effort kernels during GPU idle periods.
@@
To efficiently co-locate high-priority and best-effort tasks on a GPU with performance isolation guarantees, it is essential to schedule best-effort kernels in a controlled manner when the GPU is idle and promptly switch to high-priority kernels when needed. However, modern GPUs lack kernel scheduling primitives to terminate a running kernel and only allow context switching between kernels at the granularity of entire kernels (reef2022, ; GILMAN2021102234, ). This limitation introduces significant overhead on high-priority tasks, especially since many best-effort DL workloads involve long-running kernels. To overcome this challenge and enable finer-grained kernel scheduling, Tally uses two block-level kernel scheduling primitives: slicing and preemption. Slicing divides large kernels into smaller segments, while preemption allows for the interruption of active kernels. Our key insight is that these two scheduling primitives can be implemented in a non-intrusive and task-agnostic manner through a series of GPU kernel transformation passes, leveraging the GPU programming model which organizes kernels as collections of independent thread blocks.
@@
While slicing and preemption are effective mechanisms to manage co-execution interference, each has its trade-offs. Kernel slicing is a lightweight transformation but can incur higher overhead due to multiple kernel launches. In contrast, kernel preemption requires only a single kernel launch but involves a more complex transformation with potential synchronization overhead. Due to these differences, the optimal choice of scheduling primitives and configuration parameters (e.g., the degree of slicing) varies across different kernels. Towards this end, Tally uses a profile-guided resource provisioning strategy in its scheduling algorithm that transparently profiles the performance of best-effort kernels on-the-fly across various scheduling configurations to select the best option that meets the latency requirements of the high-priority task. This approach ensures that interference from best-effort tasks remains within acceptable limits while striving for optimal system throughput.
@@
To evaluate Tally’s ability to provide effective performance isolation, we prepare a comprehensive benchmark suite consisting of six training and six inference workloads that cover a wide range of popular DL models and frameworks (details shown in table 2). We evaluate the co-located performance of latency-critical inference tasks and best-effort training tasks by simulating the production traffic patterns for the inference tasks using the publicly available Microsoft Azure Function Trace 2021 (MAF2) dataset (maf2, ). Across all the combinations, Tally incurs a mere $7.2\%$ overhead on average in the $99^{th}$-percentile latency of inference tasks, compared to $252.3\%$, $345.0\%$, $195.5\%$, and $188.9\%$ exhibited by state-of-the-art non-intrusive GPU sharing solutions, Time-Slicing (GILMAN2021102234, ), MPS, MPS-Priority (nvidia-mps, ), and TGS (tgs2023, ) respectively. At the same time, Tally is able to achieve $80\%$ of the system throughput attained by TGS, demonstrating the effectiveness of Tally in maximizing the GPU efficiency while providing strong performance isolation guarantees.
@@
  • We introduce a novel approach to achieve fine-grained GPU kernel scheduling in a non-intrusive and task-agnostic manner using kernel slicing and preemption primitives. We further propose a priority-aware kernel scheduler that dynamically selects the appropriate scheduling primitives to optimize GPU throughput while upholding performance isolation guarantees.
  • We evaluate Tally on a diverse range of training and inference workloads and show that Tally can achieve system throughput on-par with state-of-the-art GPU sharing solutions while ensuring strong performance isolation.
@@ 2. BACKGROUND
Modern large-scale DL clusters are commonly shared among numerous users within an organization and are designed to support a diverse range of workloads (xiao2020, ; lucid, ). These include training, fine-tuning, and inference tasks for a variety of model architectures, such as convolutional neural networks (CNNs) (Qi_2017_CVPR, ; kaiming2016, ), transformer-based large language models (LLMs) (touvron2023llama, ; gpt-neo, ; devlin-etal-2019-bert, ), graph neural networks (GNNs) (graph-transformer2019, ; graph-attention2018, ), and reinforcement learning (RL) models (schulman2017proximal, ; Silver2016, ). These workloads are implemented using various DL frameworks, including widely adopted open-source platforms like PyTorch (paszke2019pytorch, ), TensorFlow (tensorflow2016, ), and JAX (jax2018github, ), as well as proprietary software stacks (Weng2022mlaas, ). Given the diversity and compute-intensive nature of DL workloads, general-purpose GPUs have become the predominant accelerator in DL clusters for their expressive programming model and massive parallel processing capabilities.
Figure 1. Overview of the GPU programming model.
@@
GPU architecture and programming model. Despite variations across vendors (e.g, NVIDIA, AMD), the core architecture and programming model of modern GPUs in DL clusters remain largely similar. GPUs are composed of hundreds of parallel processing units called Streaming Multiprocessors (SMs). The computational tasks executed on GPUs, known as kernels, are programmed using specialized languages like CUDA (cuda2024, ). As illustrated in Figure 1, GPU kernels follow a hierarchical thread group structure consisting of threads, warps, thread blocks, and grids. Threads within a kernel identify themselves using built-in variables such as threadIdx and blockIdx to determine their task assignments. Threads within a warp execute instructions in a SIMT (Single Instruction, Multiple Threads) fashion, where multiple threads execute the same instruction on different pieces of data simultaneously. All warps within a thread-block are scheduled onto a single SM, enabling them to share data using fast on-chip shared memory. Additionally, threads within a block can synchronize their execution using synchronization primitives to coordinate memory accesses. Finally, thread blocks within a grid can be scheduled and executed independently in any order, either sequentially or in parallel (cuda2024, ). This independent scheduling capability allows thread blocks to scale effectively across multiple SMs.
@@
GPU sharing systems. The diverse nature of DL workloads, coupled with their inherent inefficiencies, has led to significant concerns about GPU underutilization in many production DL clusters (hu2021characterization, ; xiao2020, ). To address this challenge, researchers have proposed numerous cluster-level optimizations (xiao2018, ; xiao2020, ; wang2021horizontally, ; salus2020, ), among which GPU sharing has emerged as a promising approach to improve utilization by consolidating multiple low-utilization jobs onto the same GPU. From an implementation perspective, current GPU sharing solutions can be broadly categorized into two levels: application-level and runtime-level approaches. Application-level solutions typically involve the introduction of new programming interfaces designed for concurrent DL execution. For example, Salus (salus2020, ) exposes a set of scheduling primitives for defining the iteration boundaries of a training job, enabling the scheduler to manage different training workloads at the iteration-level granularity. Similarly, HFTA (wang2021horizontally, ) provides APIs that enable developers to group multiple hyper-parameter tuning jobs into a single application, facilitating efficient execution on a single GPU.
@@
In contrast to application-level solutions that require modifications to user-level applications, runtime-level GPU sharing solutions operate transparently by integrating directly into either DL frameworks (xiao2020, ) (e.g., PyTorch (paszke2019pytorch, ), TensorFlow (tensorflow2016, )), container orchestration systems (tgs2023, ) (e.g., Kubernetes (kubernetes, )), or the GPU driver (reef2022, ; nvidia-mps, ; nvidia-mig, ). For example, NVIDIA GPUs provide two built-in driver-level concurrency mechanisms: Time-Slicing (GILMAN2021102234, ) and MPS (nvidia-mps, ), which enable transparent sharing of GPU resources in temporal and spatial manners respectively. Additionally, NVIDIA’s Multi-Instance GPU (MIG) (nvidia-mig, ) introduces hardware-level resource partitioning primitives that divide GPU resources, such as SMs, cache, and memory bandwidth, into multiple virtual GPU instances, each operating within an isolated environment. Besides vendor-provided solutions, TGS (tgs2023, ) integrates with Kubernetes to schedule workloads at the GPU kernel level, while REEF (reef2022, ) extends the GPU driver to offer more fine-grained thread-level scheduling capabilities. The transparent design of these solutions facilitates broader adoption across applications. However, designing these solutions is often challenging due to the lack of application-level context.